What is the maximum concentration of Ni2+ in a solution of pH 10.00? Ksp (NI(OH)2) = 2.0*10-15 M
Answer #1

The pH of the given solution is 10. pH + POH = 14 POH = 14 - pH = 14-10 pOH = -log[OH] = 4.0 -log[OH-] = 4.0 [OH]=1040 =1.0x

The equation for solubility product is written as follows: Solubility product K = [Ni?][OH-] = (sz)(1.0x10*+28) Assuming, 1.0